# Hybrid satellite-fiber quantum network

### v1

@@ X Varying the ground station telescope size
The entanglement distribution rate through the satellite is proportional to the square of the radius of the ground station telescope, as expressed in Eq. (6). Therefore, increasing the telescope radius can further improve the performance of the hybrid entanglement distribution protocol, as shown in Fig S4. We emphasize that the target end-to-end fidelity does not depend on the entanglement distribution rate (and thus on the telescope size).
@@
Figure S4: End-to-end entanglement distribution rate for ground station telescopes of increasing diameter: $a_{r}=0.5$ m (blue), $1.0$ m (purple), and $2.0$ m (red). The data are binned to facilitate comparison across cases, with each point representing an average and the shades represent the standard errors. The case $a_{r}=0.5$ m corresponds to the scattered plot in Fig. 6(i).
XI Impact of the decoherence time
@@
In the main text, we assume that the decoherence in the qubits stored in quantum memories is negligible. Here, we further investigate its impact on the end-to-end fidelity degradation. As described in Eq. (4), we model decoherence using dephasing noise. In this model, an arbitrary initial state $\rho_{0}$ evolves as
@@
where $q(t)=(1-e^{-t^{2}/\tau_{q}^{2}})/2$ and $\tau_{q}$ is the decoherence time. After entanglement is established between neighboring trapped-ion repeaters, the qubits are stored in quantum memories while waiting for the next link to complete entanglement purification. Figure S5 shows, for different values of $\tau_{q}$, the resulting end-to-end fidelity between randomly selected Alice and Bob pairs using only optical fibers as in Fig. 4(b). Shorter decoherence times lead to decreased performance of the fiber-based protocol, as expected. However, the end-to-end fidelity for $\tau_{q}$ as small as $5$ s is comparable to that for $\tau_{q}=\infty$.
Figure S5: End-to-end entanglement fidelity via optical fiber when fidelity reduction is considered for different decoherence times: $\tau_{q}=\infty$ [blue, same as in Fig. 4(b)], $\tau_{q}=5$ s (orange), and $\tau_{q}=1$ s (green).
XII Impact of the Satellite location
We assume that at all times there is at least one MEO satellite above the contiguous United States available for entanglement distribution. Because the satellite location changes in time, we evaluate the impact of satellite location on entanglement distribution efficiency. Specifically, we sample a grid of subsatellite locations for a grid spacing ${}^{\text{o}}$ in both latitude and longitude. The subsatellite locations are placed within or near the U.S. border [orange dots in Fig. S6(a)] so that every node in the network has at least one subsatellite point within a circle of diameter $\sqrt{2}\times 1^{\text{o}}$ centered on it. For each sampled location, we calculate the corresponding entanglement distribution rates and end-to-end fidelity for the same 10 0001000010\,00010 000 Alice-Bob pairs considered in Fig. 4. Figure S6(b)-(c) show the resulting distribution rates and fidelities across satellite locations via optical fibers (blue) or satellites (orange), compared with the case in which we approximate the satellite location as being at the center of the map (black). These results indicate that, as long as the satellite is positioned above the contiguous U.S., its exact location has only a small effect on the performance of the entanglement distribution protocol, supporting the validity of the satellite location approximation used in the main text.
Figure S6: End-to-end entanglement distribution rate and fidelity for different MEO satellite locations. (a) Grid of different MEO subsatellite locations (orange dots) across the contiguous U.S. and near its borders. In this representation, the longitude and latitude are put in the same scale. (b) Distribution rate through optical fibers (blue) and satellites (orange), where the latter considers the different satellite locations in (a) for each Alice-Bob pair. The black dots indicate the corresponding results from Fig. 4(a), where the satellite location is approximated to be at the center of the U.S. (c) Corresponding fidelity of the end-to-end entanglement for the time needed to establish one entanglement, where the color code is the same as in (b). The small differences between these and the results in Fig. 4 are due to the dependence of diffraction and atmospheric extinction on the satellite’s position.

### v2

@@ XI Appendix C: Varying the ground station telescope size
The entanglement distribution rate through the satellite is proportional to the square of the radius of the ground-station telescope, as expressed in Eq. (6). Therefore, increasing the telescope radius can further improve the performance of the hybrid entanglement distribution protocol, as shown in Fig. 11. We emphasize that the target end-to-end fidelity does not depend on the entanglement distribution rate (and thus on the telescope size).
@@
Figure 11: End-to-end entanglement distribution rate for ground-station telescopes of increasing diameter: $a_{\text{r}}=0.5$ m (blue), $1.0$ m (purple), and $2.0$ m (red). The data are binned to facilitate comparison across cases, with each point representing an average and the shaded bands representing the standard errors. The case $a_{\text{r}}=0.5$ m corresponds to the scattered plot in Fig. 6(i).
XII Appendix D: Impact of the decoherence time
@@
Our main results are presented assuming that the decoherence in the qubits stored in quantum memories is negligible. Here, we explicitly investigate the impact of decoherence on the end-to-end fidelity degradation. As described in Eq. (4), we model decoherence using dephasing noise. In this model, an arbitrary initial state $\rho_{0}$ evolves as
@@
where $q(t)=(1-e^{-t^{2}/\tau_{q}^{2}})/2$ and $\tau_{q}$ is the decoherence time. After entanglement is established between neighboring trapped-ion repeaters, the qubits are stored in quantum memories while waiting for the next link to complete entanglement purification. Figure 12 shows, for different values of $\tau_{q}$, the resulting end-to-end fidelity between randomly selected Alice-and-Bob pairs using only optical fibers as in Fig. 4(b). Shorter decoherence times lead to decreased performance of the fiber-based protocol, as expected. However, the end-to-end fidelity for $\tau_{q}$ as small as $5$ s is comparable to that for $\tau_{q}=\infty$.
XIII Appendix E: Impact of the Satellite location
[…]
Figure 12: End-to-end entanglement fidelity via optical fiber when fidelity reduction is considered for different decoherence times: $\tau_{q}=\infty$ [blue, same as in Fig. 4(b)], $\tau_{q}=5$ s (orange), and $\tau_{q}=1$ s (green).
Figure 13: End-to-end entanglement distribution rate and fidelity for different MEO satellite locations. (a) Grid of different MEO subsatellite locations (orange dots) across the contiguous U.S. and near its borders. (b) Distribution rate through optical fibers (blue) and satellites (orange), where the latter considers the different satellite locations in (a) for each Alice-Bob pair. The black dots indicate the corresponding results from Fig. 4(a), where the satellite location is approximated to be at the center of the U.S. (c) Corresponding fidelity of the end-to-end entanglement for the time needed to establish one entanglement, where the color code is the same as in (b). The small differences between these and the results in Fig. 4 are due to the dependence of diffraction and atmospheric extinction on the satellite’s position.